# A categorical interpretation of Morita equivalence for dynamical von Neumann algebras

### v1

@@ Abstract.
Let $\mathbb{G}$ be a locally compact quantum group and $(M,\alpha)$ a $\mathbb{G}$-$W^{*}$-algebra. The object of study of this paper is the $W^{*}$-category $\operatorname{\mathrm{Rep}}{G}(M)$ of normal, unital $\mathbb{G}$-representations of $M$ on Hilbert spaces endowed with a unitary $\mathbb{G}$-representation. This category has a right action of the category $\operatorname{\mathrm{Rep}}(\mathbb{G})=\operatorname{\mathrm{Rep}}{G}(\mathbb{C})$ for which it becomes a right $\operatorname{\mathrm{Rep}}(\mathbb{G})$-module $W^{*}$-category. Given another $\mathbb{G}$-$W^{*}$-algebra $(N,\beta)$, we denote the category of normal *-functors $\operatorname{\mathrm{Rep}}{G}(N)\to\operatorname{\mathrm{Rep}}{G}(M)$ compatible with the $\operatorname{\mathrm{Rep}}(\mathbb{G})$-module structure by $\operatorname{Fun}_{\operatorname{\mathrm{Rep}}(\mathbb{G})}(\operatorname{\mathrm{Rep}}{G}(N),\operatorname{\mathrm{Rep}}{G}(M))$ and we denote the category of $\mathbb{G}$-$M$-$N$-correspondences, studied in [DCDR23], by $\operatorname{Corr}{G}(M,N)$. We prove that there are canonical functors $P:\mathrm{Corr}{G}(M,N)\to\operatorname{Fun}_{\operatorname{\mathrm{Rep}}(\mathbb{G})}(\operatorname{\mathrm{Rep}}{G}(N),\operatorname{\mathrm{Rep}}{G}(M))$ and $Q:\operatorname{Fun}_{\operatorname{\mathrm{Rep}}(\mathbb{G})}(\operatorname{\mathrm{Rep}}{G}(N),\operatorname{\mathrm{Rep}}{G}(M))\to\operatorname{Corr}{G}(M,N)$ such that $Q\circ P\cong\operatorname{\mathrm{id}}.$ We use these functors to show that the $\mathbb{G}$-dynamical von Neumann algebras $(M,\alpha)$ and $(N,\beta)$ are equivariantly Morita equivalent if and only if $\operatorname{\mathrm{Rep}}{G}(N)$ and $\operatorname{\mathrm{Rep}}{G}(M)$ are equivalent as $\operatorname{\mathrm{Rep}}(\mathbb{G})$-module-$W^{*}$-categories. Specializing to the case where $\mathbb{G}$ is a compact quantum group, we prove that moreover $P\circ Q\cong\operatorname{\mathrm{id}}$, so that the categories $\mathrm{Corr}{G}(M,N)$ and $\operatorname{Fun}_{\operatorname{\mathrm{Rep}}(\mathbb{G})}(\operatorname{\mathrm{Rep}}{G}(N),\operatorname{\mathrm{Rep}}{G}(M))$ are equivalent. This is an equivariant version of the Eilenberg-Watts theorem for actions of compact quantum groups on von Neumann algebras.
@@ 2.1. Locally compact quantum groups
A Hopf-von Neumann algebra is a pair $(M,\Delta)$ where $M$ is a von Neumann algebra and $\Delta:M\to M\bar{\otimes}M$ a unital, normal, isometric *-homomorphism such that $(\Delta\otimes\operatorname{\mathrm{id}})\circ\Delta=(\operatorname{\mathrm{id}}\otimes\Delta)\circ\Delta$.
A locally compact quantum group $\mathbb{G}$ is a Hopf-von Neumann algebra $(L^{\infty}(\mathbb{G}),\Delta)$ for which there exist normal, semifinite, faithful weights $\Phi,\Psi:L^{\infty}(\mathbb{G})_{+}\to[0,\infty]$ such that $(\operatorname{\mathrm{id}}\otimes\Phi)\Delta(x)=\Phi(x)1$ for all $x\in\mathscr{M}_{\Phi}^{+}$ and $(\Psi\otimes\operatorname{\mathrm{id}})\Delta(x)=\Psi(x)1$ for all $x\in\mathscr{M}_{\Psi}^{+}$. The von Neumann algebra predual $L^{1}(\mathbb{G}):=L^{\infty}(\mathbb{G})_{*}$ is a completely contractive Banach algebra for the multiplication
@@
A locally compact quantum group $\mathbb{G}$ is called compact if $C_{0}(\mathbb{G})$ is unital, and we then write $C_{0}(\mathbb{G})=C(\mathbb{G})$. The canonical dense Hopf -subalgebra of $C(\mathbb{G})$ will be denoted by $\mathcal{O}(\mathbb{G})$. In that case, the left Haar weight $\Phi$ is a normal state and $\Phi=\Psi$ (after appropriate normalisations)
@@
A unitary representation of a locally compact quantum group $\mathbb{G}$ on a Hilbert space consists of a pair $(\mathcal{H},U)$ where $\mathcal{H}$ is a Hilbert space and $U\in B(\mathcal{H})\bar{\otimes}L^{\infty}(\mathbb{G})$ a unitary satisfying $(\operatorname{\mathrm{id}}\otimes\Delta)(U)=U_{12}U_{13}$. We write $(\mathcal{H},U)\in\operatorname{\mathrm{Rep}}(\mathbb{G})$. For example, we have $V\in\operatorname{\mathrm{Rep}}(\mathbb{G})$ and $W_{21}\in\operatorname{\mathrm{Rep}}(\mathbb{G})$. Given a Hilbert space $\mathcal{H}$, we write $\mathbb{I}=1\otimes 1\in B(\mathcal{H})\bar{\otimes}L^{\infty}(\mathbb{G})$ for the trivial representation. Given $(\mathcal{H},U_{\mathcal{H}}),(\mathcal{K},U_{\mathcal{K}})\in\operatorname{\mathrm{Rep}}(\mathbb{G})$, we also have that $(\mathcal{H}\otimes\mathcal{K},U_{\mathcal{H},13}U_{\mathcal{K},23})\in\operatorname{\mathrm{Rep}}(\mathbb{G})$. Given $U\in B(\mathcal{H})\bar{\otimes}L^{\infty}(\mathbb{G})$ and $\omega\in L^{1}(\mathbb{G})$, we write $U(\omega):=(\operatorname{\mathrm{id}}\otimes\omega)(U)$.
@@ 2.3. Equivariant correspondences
We call $(M,\alpha)$ and $(N,\beta)$ $\mathbb{G}$-$W^{*}$-Morita equivalent, and we write $(M,\alpha)\sim{G}(N,\beta)$, if there exists $(\mathcal{H},\pi,\rho,U)\in\operatorname{Corr}{G}(M,N)$ such that $\pi$ and $\rho$ are faithful and such that $\pi(M)^{\prime}=\rho(N).$ If $M,N$ are Morita equivalent von Neumann algebras, then we simply write $M\sim N$. This is the situation where $\mathbb{G}$ is the trivial group. It is not so hard to see that $\sim{G}$ is an equivalence relation. The notion of equivariant Morita equivalence for dynamical von Neumann algebras will be studied in depth in future work.
@@
Crossed products. Given $\mathcal{H}=(\mathcal{H},\pi,\rho,U)\in\mathrm{Corr}{G}(M,N)$, it was shown in [DCDR23, Proposition 5.20] that the following data defines a $\check{\mathbb{G}}$-correspondence $\mathcal{H}^{\rtimes}\in\mathrm{Corr}^{\check{\mathbb{G}}}(M\rtimes_{\alpha}\mathbb{G},N\rtimes_{\beta}\mathbb{G}):$
@@
The converse implication follows from the implication we just proved and Takesaki-Takai duality. ∎
@@
For more details, the reader is refered to [DCDR23, Proposition 5.5].
@@
Let us remind the reader to some important objects in $\operatorname{\mathrm{Rep}}{G}(M)$:
@@ 2.5. Generators
  • Since every $\mathbb{G}$-$M$-$N$ correspondence decomposes as a direct sum of cyclic $\mathbb{G}$-$M$-$N$-correspondences [DCDR23, Proposition 2.11], there exists a $\mathbb{G}$-$M$-$N$-correspondence $\mathcal{H}$ that contains (an isomorphic copy of) every cyclic $\mathbb{G}$-$M$-$N$-correspondence. Then $\mathcal{H}$ is a generator for $\operatorname{Corr}{G}(M,N)$, so generators for $\mathrm{Corr}{G}(M,N)$ always exist.
@@ 3. Functors associated to pairs of dynamical systems
as follows easily from [DCDR23, Proposition 5.9] and associativity of the Connes fusion tensor product.
@@ 4. An equivariant Eilenberg-Watts theorem for compact quantum groups
which is impossible since Morita equivalent von Neumann algebras have isomorphic centers [Rie74, Proposition 8.1]. Define then
@@
${}_{N}\mathscr{L}{G}(L^{2}(N)\otimes L^{2}(\mathbb{G})\otimes L^{2}(\mathbb{G}))=V_{23}^{*}U_{\alpha,13}^{*}[\rho_{N\rtimes_{\beta}\mathbb{G}}(N\rtimes_{\beta}\mathbb{G})\bar{\otimes}B(L^{2}(\mathbb{G}))]_{132}U_{\alpha,13}V_{23}.$
@@
$\begin{cases}V_{23}^{*}U_{\alpha,13}^{*}(\rho_{N}(n)\otimes 1\otimes 1)U_{\alpha,13}V_{23}&n\in N\\ 1\otimes V^{*}(1\otimes\hat{y})V&\hat{y}\in L^{\infty}(\hat{\mathbb{G}})\\ 1\otimes V^{*}(x\otimes 1)V&x\in B(L^{2}(\mathbb{G}))\end{cases}$
@@
$s_{3}F(V_{23}^{*}U_{\alpha,13}^{*}(\rho_{N}(n)\otimes 1\otimes 1)U_{\alpha,13}V_{23})=G(V_{23}^{*}U_{\alpha,13}^{*}(\rho_{N}(n)\otimes 1\otimes 1)U_{\alpha,13}V_{23})s_{3},\quad n\in N.$ (4.4)
@@
Acknowledgments: The research of the author was supported by Fonds voor Wetenschappelijk Onderzoek (Flanders), via an FWO Aspirant fellowship, grant 1162522N. The author would like to thank K. De Commer for useful input and discussions throughout this entire project. We thank L. Rollier and J. Krajczok for a useful discussion.

### v2

@@ Abstract.
Let $\mathbb{G}$ be a locally compact quantum group and $(M,\alpha)$ a $\mathbb{G}$-$W^{*}$-algebra. The object of study of this paper is the $W^{*}$-category $\operatorname{\mathrm{Rep}}{G}(M)$ of normal, unital $\mathbb{G}$-representations of $M$ on Hilbert spaces endowed with a unitary $\mathbb{G}$-representation. This category has a right action of the category $\operatorname{\mathrm{Rep}}(\mathbb{G})=\operatorname{\mathrm{Rep}}{G}(\mathbb{C})$ for which it becomes a right $\operatorname{\mathrm{Rep}}(\mathbb{G})$-module $W^{*}$-category. Given another $\mathbb{G}$-$W^{*}$-algebra $(N,\beta)$, we denote the category of normal *-functors $\operatorname{\mathrm{Rep}}{G}(N)\to\operatorname{\mathrm{Rep}}{G}(M)$ compatible with the $\operatorname{\mathrm{Rep}}(\mathbb{G})$-module structure by $\operatorname{Fun}_{\operatorname{\mathrm{Rep}}(\mathbb{G})}(\operatorname{\mathrm{Rep}}{G}(N),\operatorname{\mathrm{Rep}}{G}(M))$ and we denote the category of $\mathbb{G}$-$M$-$N$-correspondences, studied in [DCDR24], by $\operatorname{Corr}{G}(M,N)$. We prove that there are canonical functors $P:\mathrm{Corr}{G}(M,N)\to\operatorname{Fun}_{\operatorname{\mathrm{Rep}}(\mathbb{G})}(\operatorname{\mathrm{Rep}}{G}(N),\operatorname{\mathrm{Rep}}{G}(M))$ and $Q:\operatorname{Fun}_{\operatorname{\mathrm{Rep}}(\mathbb{G})}(\operatorname{\mathrm{Rep}}{G}(N),\operatorname{\mathrm{Rep}}{G}(M))\to\operatorname{Corr}{G}(M,N)$ such that $Q\circ P\cong\operatorname{\mathrm{id}}.$ We use these functors to show that the $\mathbb{G}$-dynamical von Neumann algebras $(M,\alpha)$ and $(N,\beta)$ are equivariantly Morita equivalent if and only if $\operatorname{\mathrm{Rep}}{G}(N)$ and $\operatorname{\mathrm{Rep}}{G}(M)$ are equivalent as $\operatorname{\mathrm{Rep}}(\mathbb{G})$-module-$W^{*}$-categories. Specializing to the case where $\mathbb{G}$ is a compact quantum group, we prove that moreover $P\circ Q\cong\operatorname{\mathrm{id}}$, so that the categories $\mathrm{Corr}{G}(M,N)$ and $\operatorname{Fun}_{\operatorname{\mathrm{Rep}}(\mathbb{G})}(\operatorname{\mathrm{Rep}}{G}(N),\operatorname{\mathrm{Rep}}{G}(M))$ are equivalent. This is an equivariant version of the Eilenberg-Watts theorem for actions of compact quantum groups on von Neumann algebras.
@@ 2.1. Locally compact quantum groups
A von Neumann bialgebra is a pair $(M,\Delta)$ where $M$ is a von Neumann algebra and $\Delta:M\to M\bar{\otimes}M$ a unital, normal, isometric *-homomorphism such that $(\Delta\otimes\operatorname{\mathrm{id}})\circ\Delta=(\operatorname{\mathrm{id}}\otimes\Delta)\circ\Delta$. Here, $M\bar{\otimes}M$ denotes the von Neumann algebra tensor product, containing the algebraic tensor product $M\odot M$ as a $\sigma$-weakly dense *-subalgebra.
A locally compact quantum group $\mathbb{G}$ is a von Neumann bialgebra $(L^{\infty}(\mathbb{G}),\Delta)$ for which there exist normal, semifinite, faithful weights $\Phi,\Psi:L^{\infty}(\mathbb{G})_{+}\to[0,\infty]$ such that $(\operatorname{\mathrm{id}}\otimes\Phi)\Delta(x)=\Phi(x)1$ for all $x\in\mathscr{M}_{\Phi}^{+}$ and $(\Psi\otimes\operatorname{\mathrm{id}})\Delta(x)=\Psi(x)1$ for all $x\in\mathscr{M}_{\Psi}^{+}$. The von Neumann algebra predual $L^{1}(\mathbb{G}):=L^{\infty}(\mathbb{G})_{*}$ is a completely contractive Banach algebra for the multiplication
@@
A locally compact quantum group $\mathbb{G}$ is called compact if $C_{0}(\mathbb{G})$ is unital, and we then write $C_{0}(\mathbb{G})=C(\mathbb{G})$. In that case, the left Haar weight $\Phi$ is a normal state and $\Phi=\Psi$ (after appropriate normalisations)
@@
A unitary representation of a locally compact quantum group $\mathbb{G}$ on a Hilbert space consists of a pair $(\mathcal{H},U)$ where $\mathcal{H}$ is a Hilbert space and $U\in B(\mathcal{H})\bar{\otimes}L^{\infty}(\mathbb{G})$ a unitary satisfying $(\operatorname{\mathrm{id}}\otimes\Delta)(U)=U_{12}U_{13}$. We write $(\mathcal{H},U)\in\operatorname{\mathrm{Rep}}(\mathbb{G})$. For example, we have $V\in\operatorname{\mathrm{Rep}}(\mathbb{G})$ and $W_{21}\in\operatorname{\mathrm{Rep}}(\mathbb{G})$, where $W_{21}=\sigma W$ with $\sigma:B(L^{2}(\mathbb{G}))\bar{\otimes}B(L^{2}(\mathbb{G}))\cong B(L^{2}(\mathbb{G}))\bar{\otimes}B(L^{2}(\mathbb{G}))$ the flip. Given a Hilbert space $\mathcal{H}$, we write $\mathbb{I}=1\otimes 1\in B(\mathcal{H})\bar{\otimes}L^{\infty}(\mathbb{G})$ for the trivial representation. Given $(\mathcal{H},U_{\mathcal{H}}),(\mathcal{K},U_{\mathcal{K}})\in\operatorname{\mathrm{Rep}}(\mathbb{G})$, we also have that $(\mathcal{H}\otimes\mathcal{K},U_{\mathcal{H},13}U_{\mathcal{K},23})\in\operatorname{\mathrm{Rep}}(\mathbb{G})$. Given $U\in B(\mathcal{H})\bar{\otimes}L^{\infty}(\mathbb{G})$ and $\omega\in L^{1}(\mathbb{G})$, we write $U(\omega):=(\operatorname{\mathrm{id}}\otimes\omega)(U)$.
@@ 2.3. Equivariant correspondences
We call $(M,\alpha)$ and $(N,\beta)$ $\mathbb{G}$-$W^{*}$-Morita equivalent, and we write $(M,\alpha)\sim{G}(N,\beta)$, if there exists $(\mathcal{H},\pi,\rho,U)\in\operatorname{Corr}{G}(M,N)$ such that $\pi$ and $\rho$ are faithful (= injective) and such that $\pi(M)^{\prime}=\rho(N).$ If $M,N$ are Morita equivalent von Neumann algebras, then we simply write $M\sim N$. This is the situation where $\mathbb{G}$ is the trivial group. It is not so hard to see that $\sim{G}$ is an equivalence relation. The notion of equivariant Morita equivalence for dynamical von Neumann algebras will be studied in depth in future work.
@@
Crossed products. Given $\mathcal{H}=(\mathcal{H},\pi,\rho,U)\in\mathrm{Corr}{G}(M,N)$, it was shown in [DCDR24, Proposition 5.21] that the following data defines a $\check{\mathbb{G}}$-correspondence $\mathcal{H}^{\rtimes}\in\mathrm{Corr}^{\check{\mathbb{G}}}(M\rtimes_{\alpha}\mathbb{G},N\rtimes_{\beta}\mathbb{G}):$
@@
Conversely, if $(M\rtimes_{\alpha}\mathbb{G},\operatorname{\mathrm{id}}\otimes\check{\Delta}_{r})\sim_{\check{\mathbb{G}}}(N\rtimes_{\beta}\mathbb{G},\operatorname{\mathrm{id}}\otimes\check{\Delta}_{r})$, then by the implication we just proved and Takesaki-Takai duality [DCDR24]*Theorem 5.25, we find
$M\sim_{\check{\check{\mathbb{G}}}}M\rtimes\mathbb{G}\times\check{\mathbb{G}}\sim_{\check{\check{\mathbb{G}}}}N\rtimes\mathbb{G}\rtimes\check{\mathbb{G}}\sim_{\check{\check{\mathbb{G}}}}N.$
From the Pontryagin biduality $\mathbb{G}\cong\check{\check{\mathbb{G}}}$, we then conclude that $M\sim{G}N.$
@@
For more details, the reader is refered to [DCDR24, Proposition 5.6].
@@
Let us mention some important objects in $\operatorname{\mathrm{Rep}}{G}(M)$:
@@ 2.5. Generators
  • Since every $\mathbb{G}$-$M$-$N$ correspondence decomposes as a direct sum of cyclic $\mathbb{G}$-$M$-$N$-correspondences [DCDR24, Proposition 2.11], there exists a $\mathbb{G}$-$M$-$N$-correspondence $\mathcal{H}$ that contains (an isomorphic copy of) every cyclic $\mathbb{G}$-$M$-$N$-correspondence. Then $\mathcal{H}$ is a generator for $\operatorname{Corr}{G}(M,N)$, so generators for $\mathrm{Corr}{G}(M,N)$ always exist.
@@ 3. Functors associated to pairs of dynamical systems
as follows easily from [DCDR24, Proposition 5.10] and associativity of the Connes fusion tensor product.
@@ 4. An equivariant Eilenberg-Watts theorem for compact quantum groups
Since Morita equivalent von Neumann algebras have isomorphic centers [Rie74, Proposition 8.1] and since the center of $\ell^{\infty}(\mathbbl{\Gamma})$ can be identified with $\ell^{\infty}(\operatorname{\mathrm{Irr}}(\check{\mathbbl{\Gamma}}))$, [^2] We use the standard notation $\operatorname{Irr}(\check{\mathbbl{\Gamma}})$ for a maximal set of pairwise non-isomorphic irreducible representations of the compact quantum group $\check{\mathbbl{\Gamma}}$. we see that this is impossible. Define then
@@
${}_{N}\mathscr{L}{G}(L^{2}(N)\otimes L^{2}(\mathbb{G})\otimes L^{2}(\mathbb{G}))=V_{23}^{*}U_{\beta,13}^{*}[\rho_{N\rtimes_{\beta}\mathbb{G}}(N\rtimes_{\beta}\mathbb{G})\bar{\otimes}B(L^{2}(\mathbb{G}))]_{132}U_{\beta,13}V_{23}.$
@@
$\begin{cases}V_{23}^{*}U_{\beta,13}^{*}(\rho_{N}(n)\otimes 1\otimes 1)U_{\beta,13}V_{23}&n\in N\\ 1\otimes V^{*}(1\otimes\hat{y})V&\hat{y}\in L^{\infty}(\hat{\mathbb{G}})\\ 1\otimes V^{*}(x\otimes 1)V&x\in B(L^{2}(\mathbb{G}))\end{cases}$
@@
$s_{3}F(V_{23}^{*}U_{\beta,13}^{*}(\rho_{N}(n)\otimes 1\otimes 1)U_{\beta,13}V_{23})=G(V_{23}^{*}U_{\beta,13}^{*}(\rho_{N}(n)\otimes 1\otimes 1)U_{\beta,13}V_{23})s_{3},\quad n\in N.$ (4.4)
@@
Acknowledgments: The research of the author was supported by Fonds voor Wetenschappelijk Onderzoek (Flanders), via an FWO Aspirant fellowship, grant 1162524N. The author would like to thank K. De Commer for useful input and discussions throughout this entire project. We thank J. Krajczok, L. Rollier and J. Vercruysse for a useful discussion.